# RTNinja: a generalized machine learning framework for analyzing random telegraph noise signals in nanoelectronic devices

### v1

@@ Introduction
With recent rapid advancements in data analytics and machine learning, which are beneficial across various scientific and technological domains, the analysis of RTN signals can be classified as an unsupervised learning problem of descriptive nature. Descriptive type of ML algorithms are designed to conduct exploratory data analysis and extract meaningful information (in-depth characteristics). This contrasts with commonly used predictive ML algorithms, which make predictions about future outputs based on unseen inputs after training on a set of known inputs and outputs. In the context of RTN signals, the observable data typically represents the progression of a complex signal as a function of time, which is a convolution of unknown sources, their activities, and intrinsic experimental noise. In this work, we propose a machine learning (ML) framework—RTNinja that can conduct automated data-driven analysis of RTN signals. As shown in Figure 1, an observable RTN signal is provided as the input information and RTNinja extracts the hidden information comprising of individual sources along with their corresponding activity. Compared to existing RTN analysis techniques, RTNinja has the following major advantages:
@@
Figure 1: Schematic illustration of RTN signal analysis, a descriptive type and unsupervised learning problem, using the RTNinja machine learning framework. Input to the framework is a convoluted RTN signal resulting from the activity of a hidden number of individual sources, along with intrinsic noise fluctuations. Without any predefined estimates of the underlying information, RTNinja analyzes the signal and outputs the total number of individual sources and their corresponding activity. Ideally, the output is an exact reproduction of the input sources ($N=N^{\prime}$) with their corresponding activity.
@@
Monte Carlo-based RTN Simulator
@@
RTN signals analysis exemplifies an unsupervised learning problem, where there is no availability of ground truth labels for the input data. In such scenarios, it is challenging to objectively evaluate the performance of an ML framework in identifying underlying patterns or structures in the data. Researchers frequently rely on heuristic methods, such as clustering quality metrics or visual inspection, to assess the performance of ML frameworks [31]. These methods are less reliable and can lead to subjective evaluations, unlike in supervised learning problems where the ground truth labels are available, which enables the direct comparison between ground truth and predicted output, facilitating a more straightforward and objective performance evaluation. In the context of RTN signals, it is possible to convert this unsupervised learning problem into a supervised learning scenario. It is a common practice in device research community to generate physics-based or statistical simulations of RTN signals, especially for individual defects studies, with predefined characteristics [10, 22, 26, 28, 30, 32]. In this work, we adopted a practice of building a Monte Carlo (MC) simulator that generates synthetic RTN signals, with all underlying attributes sampled from specific statistical distributions, as illustrated in Figure 2(a). The choice of these distributions aims at mimicking common research studies in the device research community, particularly those focused on understanding the impact of single defects on transistor characteristics. However, the RTNinja framework is designed to be agnostic to the underlying distributions. Figure 2(b) presents an example dataset generated by the MC-based RTN simulator, which is used to explain the RTNinja ML framework.
@@
Figure 2: Monte Carlo-based RTN Simulator. (a) Number of sources ($N$) is sampled from a Poisson distribution, where each source has three attributes: amplitude ($\Delta^{n}$), on-state activity, and off-state activity. Both state activities follow exponential distributions with the primary distribution parameters, on-state mean value ($\langle\tau_{on}^{n}\rangle$) and off-state mean value ($\langle\tau_{off}^{n}\rangle$), are sampled from a log-normal distribution. Extrinsic noise is added to the RTN signal in proportion to the overall signal magnitude. (b) The example dataset is simulated for a total time of $10^{3}$ units (e.g. seconds) at a sampling rate of 50 samples per unit. This particular signal constitutes four sources, with amplitudes within the range ($4\times 10^{-1}$$3$) and mean activities within the range ($3\times 10^{-1}$$5\times 10^{2}$).
RTNinja Framework
Overview
@@
The workflow representation of RTNinja ML framework is illustrated in Figure 3. RTNinja constitutes of two modules, namely LevelsExtractor and SourcesMapper.
@@
Figure 3: Workflow of RTNinja machine learning framework. Observable information i.e., an experimental RTN signal, is provided as input data to the framework. RTNinja analyzes this information and outputs the hidden individual RTN sources and their corresponding progression characteristics. The framework consists of two modules: LevelsExtractor and SourcesMapper. LevelsExtractor identifies discrete constant levels and builds an optimal model of the input data. SourcesMapper finds the most likely solution i.e., number of sources along with their amplitudes and activities, that explains the optimal model produced by LevelsExtractor. Each module consists of three sub-modules, with the objectives and underlying approaches of each sub-module clearly outlined.
LevelsExtractor
@@
Figures 4(a) and 4(b) illustrate the flowcharts for the Bayesian algorithm and the De-noiser, respectively. The Bayesian algorithm initiates with the formulation of a hypothesis, starting with a single constant level, which is modeled as a Gaussian distribution characterized by two parameters: mean ($\mu$) and standard deviation ($\sigma$). These parameters are initialized using the first data point and a priori estimated standard deviation value, with the prior probability set at 0.5. The next observed data point is then used to update the hypothesis through Bayes’ theorem, which in turn updates the prior probability via the likelihood function. This function evaluates how well the observed data fits the present hypothesis, and by combining it with the prior probability, the posterior probability is obtained. Depending on whether the posterior probability exceeds or falls below a predefined threshold $P_{threshold}$ (i.e., a hyperparameter), the belief is updated either by adding a new level (if below threshold) or by refining the existing hypothesis with updating the model parameters based on new evidence. This iterative approach leverages the posterior probability from each step as the prior for the next, continually improving parameter estimates and enhancing the understanding of $\mu$ and $\sigma$. Figure 4(c) demonstrates the application of the Bayesian algorithm on the example dataset, revealing that 11 levels are required to explain the full dataset, with four levels being predominant while the remaining levels have fewer data points.
The levels extracted from the Bayesian algorithm are then used to reconstruct the input signal, with a hyperparameter acting as a constraint to filter out noise. The De-noiser flowchart begins by revisiting the first two data points and assigning them to their closest Bayesian level. A ’continuity’ condition is imposed, whereby a transition between levels is considered a ’true’ RTN only if a predefined number of data points $c$ (i.e., a hyperparameter) belong to the same level after the transition. This condition serves as a qualitative measure of stochasticity, as the likelihood of a transition being an outlier is low if three data points ($c$ = 3) are observed close to the same level post-transition. It is important to note that any hidden source with faster RTN activity will also be considered noise, and such data points will be forcefully assigned to the present level as if no transition occurred. The time resolution of activity that can be extracted is limited by this condition. For example, if the measurement sampling time is 20 ms, the fastest activity that can be captured is 60 ms, given the continuity constraint is set to three data points ($c$ = 3). This iterative process is used to reconstruct the full RTN signal (Figure 4(d)).
@@
Figure 4: Working details and results of the LevelsExtractor module. (a) Flowchart of the Bayesian algorithm, which identifies the most suitable hypothesis (number of levels and their model parameters) based on observed data using Bayes’ theorem. The hypothesis is updated if the posterior probability falls below a user-defined threshold ($10^{-15}$ in this case). The conditional probability, $P(d_{t}\mid H)$, is calculated using a custom CDF-based proximity estimator, representing the likelihood of a data point belonging to its nearest level ($P=1$ if $d_{t}=\mu_{i}$). (b) Flowchart of the De-noiser, which assigns data points to their nearest levels, applying a custom continuity condition (hyperparameter set to $c$ = 3 data points) to ensure statistical significance for transitions identified as RTN signatures. (c) Bayesian algorithm extracts 11 levels, with the total weighted distribution of these levels (red) overlapping the dataset histogram (blue). (d) Input data is successfully reconstructed, with $\mu$ values of the 11 levels represented by dashed lines. (e) Standard deviation is estimated a priori using a median moving range method ($\sigma_{MMR}$). While a minority of the input data contains the RTN transition information, remaining portion corresponds to experimental noise at a constant level, leading to $\sigma_{MMR}$ estimation. Multiple feature models are built by varying the initial $\sigma$ value input to the Bayesian algorithm. BIC values are computed for all models, where the optimal $\sigma$ value and the corresponding feature model are determined by the lowest BIC value. This does not necessarily coincide with the $\sigma_{MMR}$ since fast and small-amplitude RTN sources might remain undetectable.
@@
The Bayesian algorithm, in conjunction with the De-noiser, is employed to construct a characteristic feature model of the data. The only information required a priori is the initial standard deviation, which serves as input for the Bayesian algorithm. This initial value is estimated using a median-based 2-by-2 range standard deviation estimator applied to the entire observed dataset, as illustrated in Figure 4(e). This value establishes the resolution threshold for the smallest detectable source amplitude and influences the spacing between adjacent levels. To construct the optimal model, the initial standard deviation value is systematically varied within an evenly spaced interval, ranging from $\sigma_{MMR}$ to the $80^{th}$ percentile in the 2-by-2 range-based data, generating multiple feature models. A critical consideration is ensuring that higher $\sigma$ values are sufficiently separated from potential source amplitudes to avoid masking important RTN information. Once the feature model is developed for each $\sigma$ value, the next step is to select the optimal model. This is done using the Bayesian Information Criterion (BIC), a statistical metric for model selection among a finite set of candidates. It is defined as:
@@
$\hat{L}$ represents the likelihood of the model given the data, $n$ is the number of data points, and $d$ is the number of parameters (in this case, the number of Bayesian levels). The likelihood term, $-2\log(\hat{L})$, evaluates the goodness of fit, where $y_{i}$ represents the observed data point and $\hat{y}_{i}$ denotes their corresponding reconstructed value. The complexity penalty term, $\log(n)d$, mitigates overfitting by imposing a penalty on models with a greater number of parameters. The model with lowest BIC is selected, as it provides the optimal trade-off between data fit and model complexity. The previously discussed Figures 4(c) and 4(d), depict the optimal feature model associated with the lowest BIC value, offering the most accurate representation of the data.
SourcesMapper
@@
Figure 5: (a) Flowchart illustrating the generation of multiple $N$-source sets. The output information from the LevelsExtractor module is used to extract transition probabilities between levels and all possible source amplitudes. Clusters and their representative amplitudes are identified in the normalized $P_{T}-\Delta$ space, enabling the combinatorial generation of $N$-sized $\Delta$ sets. (b) Markov transition matrix depicting the transition probabilities between the 11 Bayesian levels. Diagonal elements are masked for clarity, as they represent the high likelihood of remaining at the same level. The limited observability of transitions is evident in the broad distribution of activity rates associated with the hidden sources. (c) Result of the AP clustering algorithm applied to the normalized $P_{T}-\Delta$ space, revealing 13 clusters and their representative points (indicated with bold markers). (d) Flowchart for baseline estimation. For each candidate solution, the corresponding $2^{N}$-configuration set is generated. A range of baseline values is tested for each set, with the optimal baseline defined as the one that best matches the total distribution of Bayesian levels, determined using a maximum log-likelihood approach. (e) Example of a 16-configuration set with a specific baseline value derived from a candidate solution (4-source set), mapped onto the Bayesian levels distribution. The intersectional probability density values are used to compute the log-likelihood. Since only 11 levels are observed, some configurations may not align with any level, which does not pose an issue.
@@
The first step involves extracting meaningful features from the data generated by LevelsExtractor, as outlined in Figure 5a. The quantized data, representing level-to-level transitions, is modeled as a Markov process—a mathematical system in which transitions occur between states without memory of past states, a valid assumption for independent RTN signals. From this, a transition matrix is derived, containing the probabilities of transitions ($P_{T}$) between levels (Figure 5b). Simultaneously, the potential source amplitudes ($\Delta$) are calculated as the differences between the $\mu$ values of distinct levels. This results in a two-dimensional $P_{T}-\Delta$ space where unique groupings emerge, as illustrated in Supplementary Figure 1. These groupings reflect the distinct amplitude and activity characteristics of hidden sources. To extract an $N$-source set, the AP clustering algorithm is employed on a normalized $P_{T}-\Delta$ space to ensure equal-sized average error bars in both dimensions (Figure 5c). AP operates on the principle of ’message-passing’ between data points to identify distinct clusters [33]. It then automatically selects a data point from each cluster to act as the ’representative’ of that specific cluster. Unlike traditional clustering algorithms that require the number of clusters to be specified in advance, AP autonomously discovers both the clusters and their representatives. The results of the AP algorithm applied to the $P_{T}-\Delta$ space are shown in Figure 5c, where a total of 13 clusters are identified. Adjusting the AP hyperparameter settings led to slight variations in the number of clusters and representatives, as detailed in Supplementary Figure 1. Based on all these variants, an ensemble of one-dimensional representative potential amplitudes ($\Delta_{R}$) is prepared, without utilizing the representative $P_{T}$ information. To generate an $N$-source set, we first need to define the value of $N$. Given that the Bayesian levels are 11, the minimum number of hidden sources in the signal is 4 ($162^{4}=$ 16 state configurations). Multiple 4-source sets, referred to as candidate solutions, are then generated from the $\Delta_{R}$ ensemble, making the combinatorial number feasible (~700 in this case) for computational workload.
(a) (a)
(b) (b)
(c) (c)
Figure 6: (a) Cost function designed to evaluate candidate solutions, incorporating both static and dynamic metrics. The static metric maps each Bayesian level to its closest state configuration, quantifying the goodness of match based on proximity. The dynamic metric involves data reconstruction using these mapped state configurations, where simultaneous switching of multiple sources during a transition is treated as a violation. This process is exemplified for a 2-source system with four possible configurations, illustrating that transitions involving both sources changing their activity state are counted as violations. (b) Data reconstruction results for the optimal candidate solution, demonstrating successful signal reconstruction with only 5 violations out of a total of 1086 transitions. (c) Deconvolution of the optimal candidate solution into its four hidden sources, compared with the ground truth. The slow-activity sources (#1 and #3) are accurately extracted, with amplitude deviations within ±$\pm 0.5\%$. For the fast-activity sources (#2 and #4), the extracted amplitudes are within ±$\pm 1.7\%$, although some activity signatures are either not captured or shifted in occurrence instances. These discrepancies may arise from the de-noiser, where continuity constraints affect the quantized data, and/or from cost function constraints related to violated transitions.
@@
The next step is to determine the optimal candidate solution. This process begins by identifying the optimal baseline value for each candidate solution, ensuring that the total amplitudes of its state configurations align with the Bayesian levels. Figure 5d illustrates the flowchart for this process. Initially, a 16-configuration set is generated from a candidate solution. A range of baseline values is then added to this set, producing a group of configuration sets. The range of baseline values is chosen such that the resulting sets comprehensively cover the distribution of the 11 Bayesian levels. For each configuration set, a log-likelihood value is computed by projecting its amplitudes onto the total distribution of Bayesian levels, as demonstrated in Figure 5e for a random configuration set. The configuration set with the highest log-likelihood value determines the optimal baseline.
@@
Following this approach, each candidate solution is paired with a 16-configuration set that incorporates its corresponding optimal baseline value. To evaluate the effectiveness of a candidate solution, a custom cost function is developed, as illustrated in Figure 6(a). This cost function incorporates two metrics:
@@
The two metrics are combined into a single normalized metric relative to a predefined tolerance limit, which is a hyperparameter. In this scenario, the tolerance limits are set such that up to 2% of total transitions are allowed as violations, and a mismatch tolerance of one standard deviation is permitted. Both metrics are equally weighted, resulting in a maximum cost function value of 2. An ideal candidate solution would perfectly match all Bayesian levels with no violations, yielding a cost function value of 0. Using this cost function, all candidate solutions are evaluated, and the optimal solution is identified as the one with minimum cost function value. In this case, the optimal solution has a cost function of 1.249. Figure 6(b) illustrates the reconstructed RTN signal using this optimal candidate solution, which can be directly deconvolved into the activities of the four hidden sources (Figure 6(c)). Since the dataset used in this example is synthetic and labeled, the RTNinja framework’s result is compared with the ground truth, demonstrating successful analysis of the convoluted RTN signal.

### v2

@@ I Introduction
With recent rapid advancements in data analytics and machine learning, which are beneficial across various scientific and technological domains, the analysis of RTN signals can be classified as an unsupervised learning problem of descriptive nature. Descriptive type of ML algorithms are designed to conduct exploratory data analysis and extract meaningful information (in-depth characteristics). This contrasts with commonly used predictive ML algorithms, which make predictions about future outputs based on unseen inputs after training on a set of known inputs and outputs. In the context of RTN signals, the observable data typically represents the progression of a complex signal as a function of time, which is a convolution of unknown sources, their activities, and intrinsic experimental noise. In this work, we propose a machine learning (ML) framework—RTNinja that can conduct automated data-driven analysis of RTN signals. As shown in Figure LABEL:fig_focus, an observable RTN signal is provided as the input information and RTNinja extracts the hidden information comprising of individual sources along with their corresponding activity. Compared to existing RTN analysis techniques, RTNinja has the following major advantages:
@@
II Monte Carlo-based RTN Simulator
@@
RTN signals analysis exemplifies an unsupervised learning problem, where there is no availability of ground truth labels for the input data. In such scenarios, it is challenging to objectively evaluate the performance of an ML framework in identifying underlying patterns or structures in the data. Researchers frequently rely on heuristic methods, such as clustering quality metrics or visual inspection, to assess the performance of ML frameworks [31]. These methods are less reliable and can lead to subjective evaluations, unlike in supervised learning problems where the ground truth labels are available, which enables the direct comparison between ground truth and predicted output, facilitating a more straightforward and objective performance evaluation. In the context of RTN signals, it is possible to convert this unsupervised learning problem into a supervised learning scenario. It is a common practice in device research community to generate physics-based or statistical simulations of RTN signals, especially for individual defects studies, with predefined characteristics [10, 22, 26, 28, 30, 32]. In this work, we adopted a practice of building a Monte Carlo (MC) simulator that generates synthetic RTN signals, with all underlying attributes sampled from specific statistical distributions, as illustrated in Figure LABEL:fig_mcsim. The choice of these distributions aims at mimicking common research studies in the device research community, particularly those focused on understanding the impact of single defects on transistor characteristics. However, the RTNinja framework is designed to be agnostic to the underlying distributions. Figure LABEL:fig_mcsimexample presents an example dataset generated by the MC-based RTN simulator, which is used to explain the RTNinja ML framework.
@@
(a) (a) Schematic illustration of RTN signal analysis, a descriptive type and unsupervised learning problem, using the RTNinja machine learning framework. Input to the framework is a convoluted RTN signal resulting from the activity of a hidden number of individual sources, along with intrinsic noise fluctuations. Without any predefined estimates of the underlying information, RTNinja analyzes the signal and outputs the total number of individual sources and their corresponding activity. Ideally, the output is an exact reproduction of the input sources ($N=N^{\prime}$) with their corresponding activity. (b) Monte Carlo-based RTN Simulator. Number of sources (NN) is sampled from a Poisson distribution, where each source has three attributes: amplitude ($\Delta^{n}$), on-state activity, and off-state activity. Both state activities follow exponential distributions with the primary distribution parameters, on-state mean value ($\langle\tau_{on}^{n}\rangle$) and off-state mean value ($\langle\tau_{off}^{n}\rangle$), are sampled from a log-normal distribution. Extrinsic noise is added to the RTN signal in proportion to the overall signal magnitude. (c) The example dataset is simulated for a total time of $10^{3}$ units (e.g. seconds) at a sampling rate of 50 samples per unit. This particular signal constitutes four sources, with amplitudes within the range ($4\times 10^{-1}$33) and mean activities within the range ($3\times 10^{-1}$$5\times 10^{2}$). (d) Workflow of RTNinja machine learning framework. Observable information i.e., an experimental RTN signal, is provided as input data to the framework. RTNinja analyzes this information and outputs the hidden individual RTN sources and their corresponding progression characteristics. The framework consists of two modules: LevelsExtractor and SourcesMapper. LevelsExtractor identifies discrete constant levels and builds an optimal model of the input data. SourcesMapper finds the most likely solution i.e., number of sources along with their amplitudes and activities, that explains the optimal model produced by LevelsExtractor. Each module consists of three sub-modules, with the objectives and underlying approaches of each sub-module clearly outlined.
III RTNinja Framework
III.1 Overview
@@
The workflow representation of RTNinja ML framework is illustrated in Figure LABEL:fig_mlflow. RTNinja constitutes of two modules, namely LevelsExtractor and SourcesMapper.
@@
III.2 LevelsExtractor
@@
Figures LABEL:fig_bayesian and LABEL:fig_noisefilter illustrate the flowcharts for the Bayesian algorithm and the De-noiser, respectively. The Bayesian algorithm initiates with the formulation of a hypothesis, starting with a single constant level, which is modeled as a Gaussian distribution characterized by two parameters: mean ($\mu$) and standard deviation ($\sigma$). These parameters are initialized using the first data point and a priori estimated standard deviation value, with the prior probability set at 0.5. The next observed data point is then used to update the hypothesis through Bayes’ theorem, which in turn updates the prior probability via the likelihood function. This function evaluates how well the observed data fits the present hypothesis, and by combining it with the prior probability, the posterior probability is obtained. Depending on whether the posterior probability exceeds or falls below a predefined threshold $P_{threshold}$ (i.e., a hyperparameter), the belief is updated either by adding a new level (if below threshold) or by refining the existing hypothesis with updating the model parameters based on new evidence. This iterative approach leverages the posterior probability from each step as the prior for the next, continually improving parameter estimates and enhancing the understanding of $\mu$ and $\sigma$. Figure LABEL:fig_levels demonstrates the application of the Bayesian algorithm on the example dataset, revealing that 11 levels are required to explain the full dataset, with four levels being predominant while the remaining levels have fewer data points.
@@
(b) Working details and results of the LevelsExtractor module. (a) Flowchart of the Bayesian algorithm, which identifies the most suitable hypothesis (number of levels and their model parameters) based on observed data using Bayes’ theorem. The hypothesis is updated if the posterior probability falls below a user-defined threshold ($10^{-15}$ in this case). The conditional probability, $P(d_{t}\mid H)$, is calculated using a custom CDF-based proximity estimator, representing the likelihood of a data point belonging to its nearest level ($P=1$ if $d_{t}=\mu_{i}$). (b) Flowchart of the De-noiser, which assigns data points to their nearest levels, applying a custom continuity condition (hyperparameter set to $c$ = 3 data points) to ensure statistical significance for transitions identified as RTN signatures. (c) Bayesian algorithm extracts 11 levels, with the total weighted distribution of these levels (red) overlapping the dataset histogram (blue), and $\mu$ values of levels represented by dashed lines. (d) Input data is successfully reconstructed, with $\mu$ values of the 11 levels represented by dashed lines. (e) Standard deviation is estimated a priori using a median moving range method ($\sigma_{MMR}$). While a minority of the input data contains the RTN transition information, remaining portion corresponds to experimental noise at a constant level, leading to $\sigma_{MMR}$ estimation. Multiple feature models are built by varying the initial $\sigma$ value input to the Bayesian algorithm. BIC values are computed for all models, where the optimal $\sigma$ value and the corresponding feature model are determined by the lowest BIC value. This does not necessarily coincide with the $\sigma_{MMR}$ since fast and small-amplitude RTN sources might remain undetectable.
The levels extracted from the Bayesian algorithm are then used to reconstruct the input signal, with a hyperparameter acting as a constraint to filter out noise. The De-noiser flowchart begins by revisiting the first two data points and assigning them to their closest Bayesian level. A ’continuity’ condition is imposed, whereby a transition between levels is considered a ’true’ RTN only if a predefined number of data points $c$ (i.e., a hyperparameter) belong to the same level after the transition. This condition serves as a qualitative measure of stochasticity, as the likelihood of a transition being an outlier is low if three data points ($c$ = 3) are observed close to the same level post-transition. It is important to note that any hidden source with faster RTN activity will also be considered noise, and such data points will be forcefully assigned to the present level as if no transition occurred. The time resolution of activity that can be extracted is limited by this condition. For example, if the measurement sampling time is 20 ms, the fastest activity that can be captured is 60 ms, given the continuity constraint is set to three data points ($c$ = 3). This iterative process is used to reconstruct the full RTN signal (Figure LABEL:fig_levelsrecon).
@@
The Bayesian algorithm, in conjunction with the De-noiser, is employed to construct a characteristic feature model of the data. The only information required a priori is the initial standard deviation, which serves as input for the Bayesian algorithm. This initial value is estimated using a median-based 2-by-2 range standard deviation estimator applied to the entire observed dataset, as illustrated in Figure LABEL:fig_sigopt. This value establishes the resolution threshold for the smallest detectable source amplitude and influences the spacing between adjacent levels. To construct the optimal model, the initial standard deviation value is systematically varied within an evenly spaced interval, ranging from $\sigma_{MMR}$ to the $80^{th}$ percentile in the 2-by-2 range-based data, generating multiple feature models. A critical consideration is ensuring that higher $\sigma$ values are sufficiently separated from potential source amplitudes to avoid masking important RTN information. Once the feature model is developed for each $\sigma$ value, the next step is to select the optimal model. This is done using the Bayesian Information Criterion (BIC), a statistical metric for model selection among a finite set of candidates. It is defined as:
@@
$\text{where,}\quad\log(\hat{L})=-\frac{n}{2}\log(2\pi)-\frac{n}{2}\ln(\sigma^{2})-\frac{\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}{2\sigma^{2}}$
@@
III.3 SourcesMapper
@@
(c) (a) Flowchart illustrating the generation of multiple $N$-source sets. The output information from the LevelsExtractor module is used to extract transition probabilities between levels and all possible source amplitudes. Clusters and their representative amplitudes are identified in the normalized $P_{T}-\Delta$ space, enabling the combinatorial generation of NN-sized $\Delta$ sets. (b) Markov transition matrix depicting the transition probabilities between the 11 Bayesian levels. Diagonal elements are masked for clarity, as they represent the high likelihood of remaining at the same level. The limited observability of transitions is evident in the broad distribution of activity rates associated with the hidden sources. (c) Result of the AP clustering algorithm applied to the normalized $P_{T}-\Delta$ space, revealing 13 clusters and their representative points (indicated with bold markers). (d) Flowchart for baseline estimation. For each candidate solution, the corresponding $2^{N}$-configuration set is generated. A range of baseline values is tested for each set, with the optimal baseline defined as the one that best matches the total distribution of Bayesian levels, determined using a maximum log-likelihood approach. (e) Example of a 16-configuration set with a specific baseline value derived from a candidate solution (4-source set), mapped onto the Bayesian levels distribution. The intersectional probability density values are used to compute the log-likelihood. Since only 11 levels are observed, some configurations may not align with any level, which does not pose an issue.
@@
The first step involves extracting meaningful features from the data generated by LevelsExtractor, as outlined in Figure LABEL:fig_transclust. The quantized data, representing level-to-level transitions, is modeled as a Markov process—a mathematical system in which transitions occur between states without memory of past states, a valid assumption for independent RTN signals. From this, a transition matrix is derived, containing the probabilities of transitions ($P_{T}$) between levels (Figure LABEL:fig_transmatrix). Simultaneously, the potential source amplitudes ($\Delta$) are calculated as the differences between the $\mu$ values of distinct levels. This results in a two-dimensional $P_{T}-\Delta$ space where unique groupings emerge, as illustrated in Supplementary Figure 1. These groupings reflect the distinct amplitude and activity characteristics of hidden sources. To extract an $N$-source set, the AP clustering algorithm is employed on a normalized $P_{T}-\Delta$ space to ensure equal-sized average error bars in both dimensions (Figure LABEL:fig_clustering). AP operates on the principle of ’message-passing’ between data points to identify distinct clusters [33]. It then automatically selects a data point from each cluster to act as the ’representative’ of that specific cluster. Unlike traditional clustering algorithms that require the number of clusters to be specified in advance, AP autonomously discovers both the clusters and their representatives. The results of the AP algorithm applied to the $P_{T}-\Delta$ space are shown in Figure LABEL:fig_clustering, where a total of 13 clusters are identified. Adjusting the AP hyperparameter settings led to slight variations in the number of clusters and representatives, as detailed in Supplementary Figure 1. Based on all these variants, an ensemble of one-dimensional representative potential amplitudes ($\Delta_{R}$) is prepared, without utilizing the representative $P_{T}$ information. To generate an $N$-source set, we first need to define the value of $N$. Given that the Bayesian levels are 11, the minimum number of hidden sources in the signal is 4 ($2^{4}=16$ state configurations). Multiple 4-source sets, referred to as candidate solutions, are then generated from the $\Delta_{R}$ ensemble, making the combinatorial number feasible ($\sim$700 in this case) for computational workload.
@@
The next step is to determine the optimal candidate solution. This process begins by identifying the optimal baseline value for each candidate solution, ensuring that the total amplitudes of its state configurations align with the Bayesian levels. Figure LABEL:fig_baseline illustrates the flowchart for this process. Initially, a 16-configuration set is generated from a candidate solution. A range of baseline values is then added to this set, producing a group of configuration sets. The range of baseline values is chosen such that the resulting sets comprehensively cover the distribution of the 11 Bayesian levels. For each configuration set, a log-likelihood value is computed by projecting its amplitudes onto the total distribution of Bayesian levels, as demonstrated in Figure LABEL:fig_basemlle for a random configuration set. The configuration set with the highest log-likelihood value determines the optimal baseline.
(d) (a) Cost function designed to evaluate candidate solutions, incorporating both static and dynamic metrics. The static metric maps each Bayesian level to its closest state configuration, quantifying the goodness of match based on proximity. The dynamic metric involves data reconstruction using these mapped state configurations, where simultaneous switching of multiple sources during a transition is treated as a violation. This process is exemplified for a 2-source system with four possible configurations, illustrating that transitions involving both sources changing their activity state are counted as violations. (b) Data reconstruction results for the optimal candidate solution, demonstrating successful signal reconstruction with only 5 violations out of a total of 1086 transitions. (c) Deconvolution of the optimal candidate solution into its four hidden sources, compared with the ground truth. The slow-activity sources (#1 and #3) are accurately extracted, with amplitude deviations within $\pm 0.5\%$. For the fast-activity sources (#2 and #4), the extracted amplitudes are within $\pm 1.7\%$, although some activity signatures are either not captured or shifted in occurrence instances. These discrepancies may arise from the de-noiser, where continuity constraints affect the quantized data, and/or from cost function constraints related to violated transitions.
@@
Following this approach, each candidate solution is paired with a 16-configuration set that incorporates its corresponding optimal baseline value. To evaluate the effectiveness of a candidate solution, a custom cost function is developed, as illustrated in Figure LABEL:fig_costfn. This cost function incorporates two metrics:
@@
The two metrics are combined into a single normalized metric relative to a predefined tolerance limit, which is a hyperparameter. In this scenario, the tolerance limits are set such that up to 2% of total transitions are allowed as violations, and a mismatch tolerance of one standard deviation is permitted. Both metrics are equally weighted, resulting in a maximum cost function value of 2. An ideal candidate solution would perfectly match all Bayesian levels with no violations, yielding a cost function value of 0. Using this cost function, all candidate solutions are evaluated, and the optimal solution is identified as the one with minimum cost function value. In this case, the optimal solution has a cost function of 1.249. Figure LABEL:fig_sol illustrates the reconstructed RTN signal using this optimal candidate solution, which can be directly deconvolved into the activities of the four hidden sources (Figure LABEL:fig_solact). Since the dataset used in this example is synthetic and labeled, the RTNinja framework’s result is compared with the ground truth, demonstrating successful analysis of the convoluted RTN signal.